# The Pristine survey: XXVIII. The extremely metal-poor stream C-19 stretches over more than 100 degrees

### v1

@@
The Pristine survey: XXVIII. The extremely metal-poor stream C-19 stretches over more than 100 degrees
The discovery of the most metal-poor stream, C-19, provides us with a fossil record of a stellar structure born very soon after the Big Bang. In this work, we search for new C-19 members over the whole sky by combining two complementary stream-searching algorithms, STREAMFINDER and StarGO, and utilizing low-metallicity star samples from the Pristine survey as well as Gaia BP/RP spectro-photometric catalogues. We confirm twelve new members, spread over more than 100∘, using velocity and metallicity information from a set of spectroscopic follow-up programs that targeted a quasi-complete sample of our bright candidates (). From the updated set of stream members, we confirm that the stream is wide, with a stream width of pc, and dynamically hot, with a derived velocity dispersion of . The tension remains between these quantities and a purely baryonic scenario in which the relatively low-mass stream (even updated to a few ) stems from a globular cluster progenitor, as suggested by its chemical abundances. Some heating mechanism, such as preheating of the cluster in its own dark matter halo or through interactions with halo sub-structures appears necessary to explain the tension. The impact of binaries on the measured dispersion also remains unknown. Detailed elemental abundances of more stream members as well as multi-epoch radial velocities from spectroscopic observations are therefore crucial to fully understand the nature and past history of the most metal-poor stream of the Milky Way.
@@
Old stellar streams are fossil relics from the very early Universe. We can find them orbiting in the nearby halo, which allows us to measure motions and chemical abundances of each individual member with great precision. Detailed chemodynamical studies of streams open the window to understanding the formation of the first stellar structures as an alternative approach to direct observations, such as the images and spectra of high redshift galaxies ( e.g., GZH2, Zavala et al., 2024) using the James Webb Space Telescope (JWST).
@@
As age is a quantity that is very difficult to directly measure for most stars, the iron abundance of a star, which records the chemical enrichment (or lack thereof) of their birth place, is generally used as a crude proxy for age. Extremely metal-poor (EMP) stars have , which is less than 1/1000th of the solar abundance. These stars are born in a pristine environment at a time supernova events occurred sparsely and the interstellar medium was not well mixed ( e.g., Argast et al., 2000), roughly 1 Gyr after the big bang (). It is not hard to imagine that streams consisting of EMP stars come from stellar structures formed at the same epoch.
@@
In the Gaia era, stellar stream searches ( e.g., Ibata et al., 2021) have now revealed a handful of low-metallicity streams (Wan et al., 2020; Martin et al., 2022a), the most metal-poor of which is the C-19 stream (Martin et al., 2022b). Originally identified as an EMP stream from the Pristine photometric metallicity survey, follow-up spectroscopic observations of its brightest stars confirmed the stream stars have [Fe/H] = 3.28, with an unresolved metallicity dispersion of less than 0.18 dex at the 95 confidence level. The latter strongly suggests that the C-19 progenitor was a globular cluster. This is also supported by the relatively large variation in light element tween two members, a distinctive signature of globular clusters (see e.g., Gratton et al., 2004; Bastian & Lardo, 2018, and references within). Previous observations have also highlighted a tension between the dynamical properties and its presumed globular cluster progenitor. C-19 has an large velocity dispersion for a cluster (), with an estimated mass of only a few times , based on precise velocity measurement of 10 members brighter than G = 17.6 (Yuan et al., 2022b). This is further consolidated by the measurements of 12 faint subgiant members (G 20) by X-Shooter on ESO 8.2 m VLT (Bonifacio et al., 2024). C-19 is the most metal-poor system we have been able to obtain precise chemodynamical measurements for, and current data shows that it is a much more complicated structure than expected.
@@
Previous studies also revealed a new member star separated by 30∘ from the main body (Yuan et al., 2022b), which opens the promise that there could be more as-of-yet undiscovered C-19 members along the orbit. From the perspective of the formation of the first stellar structures, finding more bright stars in the stream would allow us to obtain their detailed elemental abundances and characterize the nature of the stream progenitor. In this work, we report a thorough search of bright C-19 members () over the whole sky. This task is now possible thanks to the newly released Gaia DR3 data and the XP spectra that have been used to derive metallicities over the full sky, combined with the improved search capabilities from the fusion of STREAMFINDER and StarGO (Yuan et al., 2018; Ibata et al., 2021; Yuan et al., 2022a). Using high-resolution spectra from telescopes in both hemispheres, we are able to confirm their memberships based on both velocity and metallicity.
@@
The identification of candidate members is explained in detail in Sec. 2. The analysis of the metallicities of the observed candidates from combined telescope programs is described in Sec. 3. The updated dynamical properties of C-19, including all the newly confirmed members, are derived in Sec. 4. The study of the chemical signature of a sub-sample of the stars reported here will be presented in a forthcoming paper (Venn et al., in preparation).
@@ 2.1 Before Gaia DR3
We first search for stars similar to known C-19 stars in dynamical space. This is done by applying StarGO to a STREAMFINDER catalog of stars likely to be in a stream, between 10 to 30 kpc. Contrary to the main Gaia EDR3 STREAMFINDER catalog (Ibata et al., 2021), we consider stars with lower probabilities, using an 8- instead of the 10- originally used by Ibata et al. (2021). StarGO uses a self-organizing map (SOM) to visualize the clustering of these stars in the (, , , ) space, where the latter two quantities characterize directions of the angular momentum vector:
@@
The 2D neuron map trained by the STREAMFINDER catalog is shown in Fig. 1(a), where a neuron is located at each grid point (, ). Stars that have similar properties in dynamical space will be located close to each other on the SOM. We can then easily select stars close to the known C-19 stars on the trained SOM. Instead of re-mapping these stars using their real measurements ( e.g., their radial velocities from the follow-up observations), we simply find them in the STREAMFINDER catalog shown as red crosses in Fig. 1(a), so they are more directly comparable to the other stars. This way, we are not biased by offsets that may arise between the actual velocities and the predicted ones from STREAMFINDER. This procedure is similar to the one we successfully applied to the search of Cetus stream members (Yuan et al., 2022a).
@@
To identify new candidate C-19 stars, we first calculate the distance of each STREAMFINDER star to the ten C-19 members on the 2D map. That is , whose distribution is shown in panel (b) of Fig. 1. We then select the 1% of stars that have the smallest . These stars are color coded by their proper motion in the direction in panels (a) and (c) of Fig. 1. Their on-sky projection (bottom panel of the Figure) immediately shows two potential stream segments (located at and ) that have the same proper motion as that of the stream orbit from Ibata et al. (2023). In total, there are nine stars in these segments brighter than , which we highlight with open yellow circles. After further excluding these candidates with photometric metallicities above in the SkyMapper DR2 (Huang et al., 2022), Pristine catalog (Starkenburg et al., 2017), we are left with four candidates along the C-19 orbit and away from the main body for spectroscopic follow-up. These are highlighted with open black circles in panel (c) of Fig. 1. Note that two of these stars are so close together, around , that they are almost indistinguishable in the Figure. We also tested tuning the threshold on to be more inclusive, but this does not lead to any more convincing candidates.
@@
We observed these four most likely candidates using the Magellan telescope on 19 August 2022 (PI: Jiang), with an exposure time of 20 - 30 minutes for each of the three targets fainter than G = 16, and 10 minutes for the most southern one (6559328209695612544) with G = 13.7. We used the standard setup of the Magellan Inamori Kyocera Echelle (MIKE) spectrograph (0.7” 5.0” slit, 22 bin, standard grating angle), which reaches a resolving power 28 000 and 35 000 in the red and blue wavelength regions, respectively (Bernstein et al., 2003). The data is reduced with the MIKE Carnegie Python pipeline (Kelson, 2003). We derive their radial velocities using iSpec (Blanco-Cuaresma et al., 2014; Blanco-Cuaresma, 2019), by comparison with standard star HD 182572 observed during the same night. All of the candidates have velocities consistent with the C-19 orbit, shown as orange circles in Fig. 2. Two stars are located close to the member separated by 30∘ away from the main body, that was previously confirmed by Yuan et al. (2022b), and the other two are located 30∘ further south. Their spectra also show that they all have metallicities 3.2, which further confirms their membership (see details in Sec. 3).
@@ 2.2 After Gaia DR3
Gaia DR3 greatly expands the Radial Velocity Sample (RVS) to with an average magnitude of (Katz et al., 2023). Moreover, using the Ca H & K band constructed from the Gaia BP/RP spectra (Carrasco et al., 2021; Andrae et al., 2023; Montegriffo et al., 2023), we are able to estimate photometric metallicities for very metal-poor stars with (the Pristine-Gaia synthetic catalog Martin et al., 2024). All this new information makes it possible to do a star-by-star search for C-19 members with the knowledge of the orbit and the metallicity of the stream.
@@
We start with stars that have Gaia RVS radial velocities, along with Pristine-Gaia synthetic metallicities (using their giant photometric metallicity model). Then we select those stars along the C-19 trajectory in the 6-D kinematic space that have differences in both tangential and radial velocities less than 30 with respect to the values predicted for the orbit. This approach gives us six potential members with . Most of them are located along the two stream segments previously found, one of which is an EMP star discovered in SEGUE (Aoki et al., 2013). There is one star that is located at , on the other side of the Milky Way disk disk.
@@
We followed up two candidates with Magellan/MIKE on August 2022 (PI: Jiang; the program already mentioned above), and one star with the Isaac Newton Telescope (INT) on 19 July 2022 (PI: Viswanathan). In the INT program, the Intermediate Dispersion Spectrograph (IDS) is equipped with the RED2 CCD. We used the R1200R grating, a 1.37” slit width, and the GG495 order-sorting filter. This setting gives a resolving power of 8000 around the CaT region from 7850 to 9150 Å(Viswanathan et al., 2024). The one star above the disk is observed with the Subaru telescope on 20 September 2022 (PI: Yuan). In this program, we used the High Dispersion Spectrograph (HDS) in the setup of StdYd with a slit width of 0.8”. This provides a wavelength coverage of 4000 – 5340 and 5450 – 6800 Å, = 45 000 (Noguchi et al., 2002). The data is reduced using the IRAF111IRAF is distributed by the National Optical Astronomy Observatory, which is operated by the Association of Universities for Research in Astronomy (AURA) under a cooperative agreement with the National Science Foundation script hdsql222http://www.subarutelescope.org/Observing/Instruments/HDS/hdsql-e.html. For each target, we took a quick exposure of 10 minutes, irrespective of the telescope. This allows us to derive their metallicities for further confirmation of their membership, which is described in detail in Sec. 3.
@@
Up to now, we are confident that, in addition to the main body of the C-19 stream, there are two additional segments of the stream at lower declination, as revealed in panel (c) of Fig. 1. The previously mentioned star-by-star search approach is based on the Gaia RVS sample, which typically contains bright stars with . We believe that there should be additional fainter members that are awaiting discovery. Based on the detailed kinematics of the confirmed members in these two stream segments, we are able to do a similar star-by-star search in fainter stars that do not have radial velocity information. To do so, we use both Pristine-Gaia synthetic metallicities from Gaia BP/RP and proprietary Pristine metallicities for fainter stars. As before, we select stars with and proper motions compatible with the orbit around the known segments. We found seven candidates with and only one of those (2641204161744171392) has a radial velocity from Gaia RVS. The velocity does agree well with expectations from the C-19 orbit at this location.
@@
We observed these seven stars with exposure time of 10 minutes each on the nights of 2 & 4 September 2023 (PI: Ibata) using the very large telescope (VLT) with the UV-visual echelle (UVES) spectrograph. The slit width was set to be 0.7”, which provides a resolving power R 60 000. The wavelength coverage is 3750 – 5000, 5700 – 7500, and 7660 – 9450 Å (Dekker et al., 2000). The observed spectra are reduced with the ESO pipeline333https://www.eso.org/sci/software/pipelines/. We confirm that all of them have radial velocities consistent with that of the C-19 orbit at their respective positions.
@@
With all these newly confirmed members, whose properties are shown in Fig. 2, the C-19 stream is continuously populated from the main body to . The other segment starts from and no members are found in between. Overall, the stream profile is more complete above due to the boundaries of the Pristine survey footprint. Within the Pristine footprint, the proprietary Pristine data has more accurate Ca HK photometry and reaches fainter stars compared to the Pristine-Gaia synthetic Ca HK from Gaia BP/RP. We therefore naturally find more fainter new members () within its sky coverage. The lack of deeper Ca HK data in the South possibly explains the gap in the current distribution of C-19. The furthest member in the North is separated from the main body of the stream by the MW disk. It is also located in a region without overlap with the Pristine footprint.
@@
2758373652717936640 | 354.57313 | 9.03539 | 14.6 | SDSS | N/A | -156.88 | 0.17 | ||
2288313499629002624† | 298.74545 | 74.65642 | 14.1 | Subaru/HDS | 600 | 48 | 23 | -134.29 | 0.67 |
2606388641446500864 | 347.06778 | -9.21263 | 17.4 | Magellan/MIKE | 1200 | 8 | 39 | -150.44 | 0.98 |
2606242131522123520 | 347.04803 | 9.42817 | 17.0 | Magellan/MIKE | 1200 | 14 | 46 | -135.14 | 0.92 |
6600784780223506944 | 339.90245 | -32.79367 | 13.5 | Magellan/MIKE | 600 | 27 | 119 | -41.15 | 0.52 |
6559328209695612544 | 327.68642 | -49.38001 | 13.7 | Magellan/MIKE | 600 | 37 | 161 | +32.42 | 0.54 |
6594796290142997376 | 335.47287 | -38.35307 | 14.5 | Magellan/MIKE | 600 | 34 | 127 | -22.09 | 0.51 |
6567859904530795904 | 331.53238 | -44.04167 | 16.5 | Magellan/MIKE | 1800 | 16 | 68 | +7.45 | 0.83 |
2641204161744171392 | 353.47032 | -0.97001 | 14.9 | VLT/UVES | 600 | 9 | 72 | -134.49 | 0.70 |
2852330082409525760 | 355.83105 | 24.77643 | 15.3 | VLT/UVES | 600 | 5 | 34 | -191.79 | 0.54 |
2760807387346283648 | 351.31022 | 7.96339 | 16.4 | VLT/UVES | 600 | 4 | 36 | -178.60 | 0.64 |
2871048855556252160 | 354.48420 | 30.22392 | 16.7 | VLT/UVES | 600 | 2 | 19 | -207.64 | 0.61 |
2658115921889849472 | 350.14296 | 2.58381 | 17.1 | VLT/UVES | 600 | 3 | 14 | -172.47 | 0.83 |
2644950674600720512 | 350.57152 | -0.22855 | 17.1 | VLT/UVES | 600 | 8 | 54 | -150.23 | 0.98 |
2640793013114783872 | 352.99372 | -1.34662 | 17.4 | VLT/UVES | 600 | 7 | 44 | -140.23 | 0.70 |

### v4

# The Pristine survey – XXVII. The extremely metal-poor stream C-19 stretches over more than 100 degrees

@@
The Pristine survey – XXVII. The extremely metal-poor stream C-19 stretches over more than 100 degrees
The discovery of the most metal-poor stream, C-19, provides us with a fossil record of a stellar structure born very soon after the big bang. In this work, we search for new C-19 members throughout the sky by combining two complementary stream-searching algorithms, STREAMFINDER and StarGO, and utilizing low-metallicity star samples from the Pristine survey, as well as Gaia BP and RP spectrophotometric catalogs. We confirm 12 new members, spread over more than 100∘, using velocity and metallicity information from a set of spectroscopic follow-up programs that targeted a quasi-complete sample of our bright candidates (). From the updated set of stream members, we confirm that the stream is wide, with a stream width of pc, and dynamically hot, with a derived velocity dispersion of . The tension remains between these quantities and a purely baryonic scenario in which the relatively low-mass stream (even updated to a few ) stems from a globular cluster progenitor, as suggested by its chemical abundances. Some heating mechanism, such as preheating of the cluster in its own dark matter halo or through interactions with halo substructures, appears necessary to explain the tension. The impact of binaries on the measured dispersion also remains unknown. Detailed elemental abundances of more stream members, as well as multi-epoch radial velocities from spectroscopic observations, are therefore crucial to fully understanding the nature and past history of the most metal-poor stream of the Milky Way.
@@
Old stellar streams are fossil relics from the very early Universe. We can find them orbiting in the nearby halo, which allows us to measure the motions and chemical abundances of each individual member with great precision. Detailed chemodynamical studies of streams open the window to understanding the formation of the first stellar structures, as an alternative approach to direct observations such as the images and spectra of high redshift galaxies ( e.g., GZH2, Zavala et al., 2024) using the James Webb Space Telescope (JWST).
@@
As age is a quantity that is very difficult to directly measure for most stars, the iron abundance of a star, which records the chemical enrichment (or lack thereof) of its birth place, is generally used as a crude proxy for age. Extremely metal-poor (EMP) stars have , which is less than 1/1000th of the solar abundance. These stars were born in a pristine environment at a time when supernova events occurred sparsely and the interstellar medium was not well mixed ( e.g., Argast et al., 2000), roughly 1 Gyr after the big bang (). It is not hard to imagine that streams consisting of EMP stars come from stellar structures formed at the same epoch.
@@
In the Gaia era, stellar stream searches ( e.g., Ibata et al., 2021) have now revealed a handful of low-metallicity streams (Wan et al., 2020; Martin et al., 2022a), the most metal-poor of which is the C-19 stream (Martin et al., 2022b). Originally identified as an EMP stream from the Pristine photometric metallicity survey, follow-up spectroscopic observations of its brightest stars confirmed the stream stars have [Fe/H] = 3.28, with an unresolved metallicity dispersion of less than 0.18 dex at the 95 confidence level. The latter strongly suggests that the C-19 progenitor was a globular cluster. This is also supported by the relatively large variation in light elements between two members, a distinctive signature of globular clusters (see, e.g., Gratton et al., 2004; Bastian & Lardo, 2018, and references within). Previous observations have also highlighted a tension between the dynamical properties and its presumed globular cluster progenitor. C-19 has a large velocity dispersion for a cluster (), with an estimated mass of only a few times , based on precise velocity measurement of ten members brighter than G = 17.6 (Yuan et al., 2022b). This is further consolidated by the measurements of 12 faint subgiant members (G 20) by X-Shooter on ESO 8.2 m Very Large Telescope (VLT) (Bonifacio et al., 2024). C-19 is the most metal-poor system we have been able to obtain precise chemodynamical measurements for, and current data show that it is a much more complicated structure than expected.
@@
Previous studies also revealed a new member star separated by 30∘ from the main body (Yuan et al., 2022b), which opens the promise that there could be more as-yet undiscovered C-19 members along the orbit. From the perspective of the formation of the first stellar structures, finding more bright stars in the stream would enable us to obtain their detailed elemental abundances and characterize the nature of the stream progenitor. In this work, we report a thorough search for bright C-19 members () throughout the sky. This task is now possible thanks to the new Gaia DR3 and the BP and RP (BP/RP) spectra that have been used to derive metallicities over the full sky, combined with the improved search capabilities from the fusion of STREAMFINDER and StarGO (Yuan et al., 2018; Ibata et al., 2021; Yuan et al., 2022a). Using high-resolution spectra from telescopes in both hemispheres, we are able to confirm their memberships based on both velocity and metallicity.
@@
The identification of candidate members is explained in detail in Sec. 2. The analysis of the metallicities of the observed candidates from combined telescope programs is described in Sec. 3. The updated dynamical properties of C-19, including all the newly confirmed members, are derived in Sec. 4. The study of the chemical signature of a subsample of the stars reported here will be presented in a forthcoming paper (Venn et al., in preparation).
@@ 2.1 Before Gaia DR3
We first searched for stars similar to known C-19 stars in dynamical space. This was done by applying StarGO to a STREAMFINDER catalog of stars likely to be in a stream, between 10 to 30 kpc. Contrary to the main Gaia EDR3 STREAMFINDER catalog (Ibata et al., 2021), we considered stars with lower probabilities, using an 8- instead of the 10- originally used by Ibata et al. (2021). StarGO uses a self-organizing map (SOM) to visualize the clustering of these stars in the (, , , ) space, where the latter two quantities characterize directions of the angular momentum vector:
@@
The 2D neuron map trained by the STREAMFINDER catalog is shown in Fig. 1(a), where a neuron is located at each grid point (, ). Stars that have similar properties in dynamical space will be located close to each other on the SOM. This enabled us to easily select stars close to the known C-19 stars on the trained SOM. Instead of remapping these stars using their real measurements ( e.g., their radial velocities from the follow-up observations), we simply found them in the STREAMFINDER catalog, shown as red crosses in Fig. 1(a), so they were more directly comparable to the other stars. In this way, we were not biased by offsets that might arise between the actual velocities and the predicted ones from STREAMFINDER. This procedure is similar to the one we successfully applied to the search for Cetus stream members (Yuan et al., 2022a).
@@
To identify new candidate C-19 stars, we first calculated the distance, , of each STREAMFINDER star to the ten C-19 members on the 2D map. This is , whose distribution is shown in panel (b) of Fig. 1. We then selected the 1% of stars that have the smallest . These stars are color coded by their proper motion in the direction in panels (a) and (c) of Fig. 1. Their on-sky projection (bottom panel of the Figure) immediately shows two potential stream segments (located at and ) that have the same proper motion as that of the stream orbit from Ibata et al. (2023). In total, there are nine stars in these segments brighter than , which we highlight with open yellow circles. After further excluding these candidates with photometric metallicities above in the SkyMapper DR2 (Huang et al., 2022) and the Pristine catalog (Starkenburg et al., 2017), we are left with four candidates along the C-19 orbit and away from the main body for spectroscopic follow-up. These are highlighted with open black circles in panel (c) of Fig. 1. We note that two of these stars are so close together, around , that they are almost indistinguishable in the Figure. We also tested tuning the threshold on to be more inclusive, but this does not lead to any candidates that are more convincing.
@@
We observed these four most likely candidates using the Magellan telescope on 19 August 2022 (PI: Jiang), with an exposure time of 20 – 30 minutes for each of the three targets fainter than G = 16.0, and 10 minutes for the most southern one (6559328209695612544) with G = 13.7. We used the standard setup of the Magellan Inamori Kyocera Echelle (MIKE) spectrograph (0.7” 5.0” slit, 22 bin, standard grating angle), which reaches a resolving power of 28 000 and 35 000 in the red and blue wavelength regions, respectively (Bernstein et al., 2003). The data were reduced with the MIKE Carnegie Python pipeline (Kelson, 2003). We derived their radial velocities using iSpec (Blanco-Cuaresma et al., 2014; Blanco-Cuaresma, 2019), by comparison with standard star HD 182572 observed during the same night. All of the candidates have velocities consistent with the C-19 orbit, listed as stars1 in Tab.4. From experience using MIKE, we note that it is very difficult to calibrate velocities extracted from its spectra at better than 2 km s-1. We therefore assumed this value as the systematic floor to the MIKE velocity uncertainties, and the statistical uncertainties are denoted as with a star symbol in the upper script in the last column of Tab.4. Given the large velocity dispersion of C-19, this choice does impact our inference on the velocity dispersion or the membership. We combined both systematic and statistical uncertainties to derive the velocity dispersion of the stream in Sec.4. These four stars are plotted as orange circles in Fig. 2. Two of them are located close to the member separated by 30∘ away from the main body, which was previously confirmed by Yuan et al. (2022b), and the other two are located 30∘ further south. Their spectra also show that they all have metallicities 3.2, which further confirms their membership (see details in Sec. 3).
@@ 2.2 After Gaia DR3
Gaia DR3 greatly expands the radial velocity sample (RVS) to , with an average magnitude of (Katz et al., 2023). Moreover, using the Ca H&K band constructed from the Gaia BP/RP spectra (Carrasco et al., 2021; Andrae et al., 2023; Montegriffo et al., 2023), we are able to estimate photometric metallicities for very metal-poor stars with (the Pristine-Gaia synthetic catalog; Martin et al., 2024). All this new information makes it possible to conduct a star-by-star search for C-19 members, using our knowledge of the orbit and the metallicity of the stream.
@@
We started with stars that have Gaia RVS radial velocities along with Pristine-Gaia synthetic metallicities (using their giant photometric metallicity model). Then we selected those stars along the C-19 trajectory in the 6D kinematic space that have differences in both tangential and radial velocities of less than 30 with respect to the values predicted for the orbit. This approach gives us five potential members with . Among them, one is an EMP star discovered in SEGUE (Aoki et al., 2013) with consistent radial velocity; therefore, it is immediately confirmed. The other candidates are located along the two stream segments previously found, except for one star that is located at , on the other side of the Milky Way disk.
@@
We followed up two candidates with Magellan/MIKE in August 2022 (PI: Jiang; the program already mentioned above) and one star with the Isaac Newton Telescope (INT) on 19 July 2022 (PI: Viswanathan). In the INT program, the intermediate dispersion spectrograph (IDS) is equipped with the RED2 CCD. We used the R1200R grating, a 1.37” slit width, and the GG495 order-sorting filter. This setting gives a resolving power of 8000 around the Ca triplet region from 7850 to 9150 Å (viswanathan2025). The star above the disk was observed with the Subaru telescope on 20 September 2022 (PI: Yuan). In this program, we used the High Dispersion Spectrograph (HDS) in the setup of StdYd, with a slit width of 0.8”. This provides a wavelength coverage of 4000 – 5340 and 5450 – 6800 Å, = 45 000 (Noguchi et al., 2002). The data were reduced using the IRAF111IRAF is distributed by the National Optical Astronomy Observatory, which is operated by the Association of Universities for Research in Astronomy (AURA) under a cooperative agreement with the National Science Foundation script hdsql222http://www.subarutelescope.org/Observing/Instruments/HDS/hdsql-e.html. For each target, we took a quick exposure of 10 minutes, irrespective of the telescope. This allowed us to derive their metallicities for further confirmation of their membership, which is described in detail in Sec. 3. These five stars are listed as stars2 in Tab.4.
@@
Up to now, we are confident that, in addition to the main body of the C-19 stream, there are two additional segments of the stream at lower declination, as revealed in panel (c) of Fig. 1. The previously mentioned star-by-star search approach is based on the Gaia RVS sample, which typically contains bright stars with . We believe that there should be additional fainter members that are awaiting discovery. Based on the detailed kinematics of the confirmed members in these two stream segments, we were able to do a similar star-by-star search in fainter stars that do not have radial velocity information. To do so, we used both Pristine-Gaia synthetic metallicities from Gaia BP/RP and proprietary Pristine metallicities for fainter stars. As before, we selected stars with and proper motions compatible with the orbit around the known segments. We find seven candidates with , only one of which (2641204161744171392) has a radial velocity from Gaia RVS. The velocity agrees well with expectations from the C-19 orbit at this location.
@@
We observed these seven stars with an exposure time of 10 minutes each on the nights of 2 and 4 September 2023 (PI: Ibata) using the Very Large Telescope (VLT) with the UV-visual echelle (UVES) spectrograph. The slit width was set to 0.7”, which provided a resolving power R 60 000. The wavelength coverage was 3750 – 5000, 5700 – 7500, and 7660 – 9450 Å (Dekker et al., 2000). The observed spectra were reduced with the ESO pipeline333https://www.eso.org/sci/software/pipelines/. We confirm that all of them have radial velocities consistent with those of the C-19 orbit at their respective positions, listed as stars3 in Tab.4. We note that there is one star⋇ (2658115921889849472) that has a quite low quality spectrum and has a difference of 6 in radial velocity derived from two bands around 5800 – 6800Å and 8400 - 8750Å. The membership is still valid despite the velocity discrepancy. But the spectrum does not have sufficient information to determine its accurate velocity as well as its metallicity. In fact, the metallicity derived from its Ca triplet region in Sec.3 is above [Fe/H]=3. Therefore, we do not consider this star as a confirmed member and do not take it into account in the analysis of velocity dispersion in Sec. 4.
@@
With all these candidates with confirmed velocities, whose properties are shown in Fig. 2, the C-19 stream is continuously populated from the main body to . The other segment starts from and no members are found in between. Overall, the stream profile is more complete above , due to the boundaries of the Pristine survey footprint. Within the Pristine footprint, the proprietary Pristine data have more accurate Ca HK photometry and reach fainter stars compared to the Pristine-Gaia synthetic Ca HK from Gaia BP/RP. We therefore naturally find more fainter new members () within its sky coverage. The lack of deeper Ca HK data in the south possibly explains the gap in the current distribution of C-19. The furthest member in the north is separated from the main body of the stream by the Milky Way disk. It is also located in a region without overlap with the Pristine footprint.
@@
| 28208669730645044482 | 353.39287 | 17.75368 | 13.8 | INT/IDS | 600 | 44 | -190.08 | 2.44 | |
| 27583736527179366402 | 354.57313 | 9.03539 | 14.6 | SDSS | N/A | -156.88 | 0.17 | ||
| 22883134996290026242† | 298.74545 | 74.65642 | 14.1 | Subaru/HDS | 600 | 48 | 23 | -134.29 | 0.67 |
| 26063886414465008641 | 347.06778 | -9.21263 | 17.4 | Magellan/MIKE | 1200 | 8 | 39 | -150.44 | 0.98∗ |
| 26062421315221235201 | 347.04803 | -9.42817 | 17.0 | Magellan/MIKE | 1200 | 14 | 46 | -135.14 | 0.92∗ |
| 66007847802235069442 | 339.90245 | -32.79367 | 13.5 | Magellan/MIKE | 600 | 27 | 119 | -41.15 | 0.52∗ |
| 65593282096956125441 | 327.68642 | -49.38001 | 13.7 | Magellan/MIKE | 600 | 37 | 161 | +32.42 | 0.54∗ |
| 65947962901429973762 | 335.47287 | -38.35307 | 14.5 | Magellan/MIKE | 600 | 34 | 127 | -22.09 | 0.51∗ |
| 65678599045307959041 | 331.53238 | -44.04167 | 16.5 | Magellan/MIKE | 1800 | 16 | 68 | +7.45 | 0.83∗ |
| 26412041617441713923 | 353.47032 | -0.97001 | 14.9 | VLT/UVES | 600 | 9 | 72 | -134.49 | 0.70 |
| 28523300824095257603 | 355.83105 | 24.77643 | 15.3 | VLT/UVES | 600 | 5 | 34 | -191.79 | 0.54 |
| 27608073873462836483 | 351.31022 | 7.96339 | 16.4 | VLT/UVES | 600 | 4 | 36 | -178.60 | 0.64 |
| 28710488555562521603 | 354.48420 | 30.22392 | 16.7 | VLT/UVES | 600 | 2 | 19 | -207.64 | 0.61 |
| 26581159218898494723⋇ | 350.14296 | 2.58381 | 17.1 | VLT/UVES | 600 | 3 | 14 | -172.47 | 0.83 |
| 26449506746007205123 | 350.57152 | -0.22855 | 17.1 | VLT/UVES | 600 | 8 | 54 | -150.23 | 0.98 |
| 26407930131147838723 | 352.99372 | -1.34662 | 17.4 | VLT/UVES | 600 | 7 | 44 | -140.23 | 0.70 |